### v2

@@ 2.2 Nonconvex-concave Setting
In this subsection, we propose a completely parameter-free AGP algorithm for solving nonconvex-concave minimax problems. Similar to the idea of the PF-AGP-NSC Algorithm, we estimate the Lipschitz constants $L_{11}$, $L_{12}$ and $L_{22}$ iteratively, however we do not need to estimate the strongly concave constant $\mu$. At each iteration of the proposed algorithm, we aim to find a tuple $(l^{k,i}_{11},l^{k,i}_{12},l^{k,i}_{22})$ by backtracking such that conditions (C1), (C2), (C3) and the following condition (C5) are satisfied at $(x_{k,i},y_{k,i})$:
$C_{5}^{k,i}$ $=\langle\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k}),y_{k,i}-y_{k}\rangle+\frac{c^{2}_{k,i}+c_{k,i}{l}^{k,i}_{22}}{{l}^{k,i}_{22}+2c_{k,i}}\|y_{k,i}-y_{k}\|^{2}$
$\qquad+\frac{1}{{l}^{k,i}_{22}+2c_{k,i}}\|\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k})\|^{2}\leq 0.$ (C5)
The proposed algorithm, denoted as PF-AGP-NC, is formally presented in Algorithm 2.
Please note that the existing SGDA-B algorithm (Xu et al., 2023) for solving nonconvex-concave minimax problems requires the realization of the diameter of the feasible set $\mathcal{Y}$ in order to construct the key auxiliary problem, while the PF-AGP-NC algorithm we propose is a completely parameter-free algorithm that does not require any prior parameters.
@@
  Step 1: Input $x_{1},y_{1}$, $\beta_{1}$, $\gamma_{1}$, $c_{1}$, $l_{11}^{1}$, $l_{12}^{1}$, $l_{22}^{1}$; Set $k=1$.
@@
   (a): Set $i=1$, $l_{11}^{k,i}=l_{11}^{k}$, $l_{12}^{k,i}=l_{12}^{k}$, $l_{22}^{k,i}=l_{22}^{k}$, $\beta_{k,i}=\beta_{k}$, $\gamma_{k,i}=\gamma_{k}$, $c_{k,i}=c_{k}$.
@@
$l^{k,i+1}_{22}$ $=\frac{\max\{\mbox{sgn}(C_{3}^{k,i}),\mbox{sgn}(C_{5}^{k,i})\}+3}{2}l^{k,i}_{22}.$ (45)
   (e): If $C_{1}^{k,i}\leq 0$, $C_{2}^{k,i}\leq 0$, $C_{3}^{k,i}\leq 0$ and $C_{5}^{k,i}\leq 0$, then
     $x_{k+1}=x_{k,i}$, $y_{k+1}=y_{k,i}$, $l_{11}^{k}=l_{11}^{k,i+1}$, $l_{12}^{k}=l_{12}^{k,i+1}$, $l_{22}^{k}=l_{22}^{k,i+1}$, $\beta_{k}=\beta_{k,i}$, $\gamma_{k}=\gamma_{k,i}$,
     $c_{k}=c_{k,i}$, go to Step 3;
@@
$\beta_{k,i}=\frac{l_{12}^{k,i}}{20l^{k-1,i}_{22}}+\frac{1280(l_{12}^{k,i})^{2}\sqrt{k}}{19^{2}l_{12}^{k-1,i}},\quad\gamma_{k,i}=20l^{k,i}_{22},\quad c_{k,i}=\frac{19l^{k,i}_{22}}{k^{1/4}},$ (46)
@@ 2.2.1 Complexity Analysis
In this subsection, we analyze the iteration complexity of Algorithm 2 for solving (P) under the general nonconvex-concave setting. The backtracking strategy (C5) is a key inequality that we will use to establish some important recursions for the PF-AGP-NC algorithm under the nonconvex-concave setting in the following two lemmas. This is also one of the main differences between proofs in the nonconvex- strongly concave setting and proofs in the nonconvex-concave setting.
@@
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. If $\forall k$, $\beta_{k}>l^{k}_{11}$ and $\gamma_{k}\geq\frac{{l}_{22}^{k}+2c_{k}}{2}$, then
@@
$f_{k}(x_{k+1},y_{k+1})-f_{k}(x_{k+1},y_{k})$ $\leq\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
@@
$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
$\quad+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$
$\quad+\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle,$ (49)
@@
Thirdly, by (C4) we have that
@@
$\leq$ $-\frac{1}{{l}_{22}^{k-1}+2c_{k-1}}\|\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1})\|^{2}$
$\quad-\frac{c_{k-1}{l}_{22}^{k-1}+c^{2}_{k-1}}{{l}_{22}^{k-1}+2c_{k-1}}\|y_{k}-y_{k-1}\|^{2}$
$\leq$ $-\frac{1}{{l}_{22}^{k-1}+2c_{k-1}}\|\nabla_{y}f_{k}(x_{k},y_{k})-\nabla_{y}f_{k}(x_{k},y_{k-1})\|^{2}.$ (52)
@@
$\begin{split}\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle=\frac{\gamma_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}-\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.\end{split}$ (53)
By $\gamma_{k}\geq\frac{{l}_{22}^{k}+2c_{k}}{2}$, plugging (50)-(53) into (49), and using the definition of $f_{k}(x_{k+1},y_{k+1})$ and $f_{k}(x_{k+1},y_{k})$, we obtain
@@
$\beta_{k}>l^{k}_{11},\quad\frac{\gamma_{k+1}}{c_{k+1}}-\frac{\gamma_{k}}{c_{k}}\leq\frac{1}{5},\quad\gamma_{k}\geq\frac{{l}^{k}_{22}+2c_{k}}{2},$ (58)
@@
$F_{k+1}-F_{k}=$ $-\left(\frac{9\gamma_{k}}{10}-\frac{c_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}$
$-\left(+\frac{\beta_{k}}{2}-\frac{(l^{k}_{12})^{2}}{2\gamma_{k-1}}-\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{c_{k-1}^{2}}\right)\|x_{k+1}-x_{k}\|^{2}$
@@
$\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1})-\nabla_{y}f_{k}(x_{k+1},y_{k})+$ $\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle\leq 0,$
@@
$=$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
@@
Using an argument similar to the proof of (50)-(53), by (C5) and the Cauchy-Schwarz inequality, we conclude from the above inequality that
@@
$-\frac{c_{k-1}{l}_{22}^{k-1}+c^{2}_{k-1}}{{l}_{22}^{k-1}+2c_{k-1}}\|y_{k}-y_{k-1}\|^{2},$ (63)
for any $a_{k}>0$. Obviously, we have $-\frac{c_{k-1}{l}^{k-1}_{22}+c^{2}_{k-1}}{2c_{k-1}+{l}^{k-1}_{22}}\leq-\frac{c_{k-1}{l}^{k-1}_{22}+c^{2}_{k-1}}{2({l}^{k-1}_{22}+c_{k-1})}=-\frac{c_{k-1}}{2}.$ Combining $\gamma_{k}\geq\frac{{l}^{k}_{22}+2c_{k}}{2}$ and rearranging the terms in (63), we obtain
@@
$\quad\leq\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}+\left(\frac{\alpha_{k}}{2}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k}\|^{2}$
$\quad\quad+\frac{(l^{k}_{12})^{2}}{2a_{k}}\|x_{k+1}-x_{k}\|^{2}-\frac{c_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}.$ (64)
@@
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. If $\beta_{k}=\frac{l_{12}^{k}}{20l^{k-1}_{22}}+\frac{1280(l_{12}^{k})^{2}\sqrt{k}}{19^{2}l_{12}^{k-1}}$, $\gamma_{k}=20l^{k}_{22}$ , $c_{k}=\frac{19l^{k}_{22}}{k^{1/4}}$, then for any given $\varepsilon>0$, $k\geq 2$, we have
$T(\varepsilon)\leq\max\left(\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$
where $d_{2}=F_{2}-\underline{F}+339L_{22}\hat{\mathcal{D}}_{y}^{2}$, $d_{3}=\max\{\bar{d}_{1},\frac{20\cdot 19^{2}L_{22}^{2}}{17(l_{12}^{1})^{2}}\}$, $\hat{\mathcal{D}}_{y}=\max\{\|y\|\,|y\in Y\}$ with $\underline{F}={\underline{f}}-639L_{22}\hat{\mathcal{D}}_{y}^{2}$, $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\bar{d}_{1}=32+\frac{1+1600L^{2}_{22}}{128({l}_{12}^{1})^{2}}$.
@@
This observation, in view of Lemma 8, $\frac{c_{k}}{2}\leq\frac{19\gamma_{k}}{40k^{1/4}}\leq\frac{19\gamma_{k}}{40}$ implies that
$\begin{split}&{\alpha_{k}}\left\|x_{k+1}-x_{k}\right\|^{2}+\frac{17\gamma_{k}}{40}\left\|y_{k+1}-y_{k}\right\|^{2}\\ \leq&F_{k}-F_{k+1}+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}.\end{split}$ (67)
@@
$\leq 32+\frac{1+1600L^{2}_{22}}{128({l}_{12}^{1})^{2}}=\bar{d}_{1}.$ (72)
@@
Denote $d_{k}^{(2)}=\frac{1}{\max\{\bar{d}_{1}\alpha_{k},{\frac{17}{80}\gamma_{k}}\}}$, $\tilde{c}_{k}=\frac{19}{20k^{1/4}}$, by multiplying $d^{(2)}_{k}$ on the both sides of (73), and using (67), we have
@@
$\sum_{k=2}^{\tilde{T}{(\varepsilon)}}d^{(2)}_{k}\left\|\nabla\tilde{G}_{k}\right\|^{2}$ $\leq F_{2}-F_{\tilde{T}(\varepsilon)}+8\left(\frac{\gamma_{\tilde{T}(\varepsilon)-1}c_{\tilde{T}(\varepsilon)}}{c_{\tilde{T}(\varepsilon)-1}}-\frac{\gamma_{1}c_{2}}{c_{1}}\right)\hat{\mathcal{D}}_{y}^{2}+40L_{22}\left(\frac{\tilde{c}_{1}}{2}-\frac{\tilde{c}_{\tilde{T}(\varepsilon)}}{2}\right)\hat{\mathcal{D}}^{2}_{y}$
@@
We can see from the selection of $d_{3}$ that $d_{3}\geq\max\{\bar{d}_{1},\frac{80\gamma_{k}}{17\alpha_{k}}\}$, which implies $d^{(2)}_{k}\geq\frac{1}{d_{3}\alpha_{k}}$, by multiplying $d_{3}$ on the both sides of (78), and combining the definition of $d_{2}$, we have $\sum_{k=2}^{\infty}\frac{1}{\alpha_{k}}\left\|\nabla\tilde{G}_{k}\right\|^{2}\leq d_{2}d_{3}$, which, by the definition of $\tilde{T}(\varepsilon)$, implies that
@@
By using the fact $\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\sqrt{k}}\geq\sqrt{\tilde{T}{(\varepsilon)}}-1$ and (79), we conclude $\frac{\varepsilon^{2}}{4}\leq\frac{1280\cdot\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\left(\sqrt{\tilde{T}(\varepsilon)}-1\right)}$ or equivalently,
$\tilde{T}(\varepsilon)\leq\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2}.$ (80)
@@
$\leq\max\left(\left(\frac{5120\left(L_{12}\right)^{2}d_{2}d_{3}}{19^{2}l_{22}^{1}\varepsilon^{2}}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$ (81)

### v3

@@ 2.2 Nonconvex-concave Setting
  Step 1: Input $x_{1},y_{1}$, $\beta_{0}$, $\gamma_{0}$, $c_{0}$, $l_{11}^{0}$, $l_{12}^{0}$, $l_{22}^{0}$; Set $k=1$.
@@
   (a): Set $i=0$$l_{11}^{k,i}={l_{11}^{k-1}}$, $l_{12}^{k,i}={l_{12}^{k-1}}$, $l_{22}^{k,i}=l_{22}^{k-1}$, $\beta_{k,i}=\beta_{k-1}$, $\gamma_{k,i}=\gamma_{k-1}$, $c_{k,i}=c_{k-1}$.
@@
$l^{k,i+1}_{22}$ $=\frac{\mbox{sgn}(C_{5}^{k,i})+3}{2}l^{k,i}_{22}.$ (45)
   (e):  If $C_{1}^{k,i}\leq 0$, $C_{2}^{k,i}\leq 0$ and $C_{5}^{k,i}\leq 0$, then
     $x_{k+1}=x_{k,i}$, $y_{k+1}=y_{k,i}$, $l_{11}^{k}=l_{11}^{k,i+1}$, $l_{12}^{k}=l_{12}^{k,i+1}$, $l_{22}^{k}=l_{22}^{k,i+1}$, $\beta_{k}=\beta_{k,i}$,
     $\gamma_{k}=\gamma_{k,i}$, $c_{k}=c_{k,i}$, go to Step 3;
@@
$\beta_{k,i}=\frac{l_{12}^{k,i}}{20l^{k-1,i}_{22}}+\frac{2(l_{12}^{k,i})^{2}\sqrt{k}}{l_{12}^{k-1,i}},\quad\gamma_{k,i}=20l^{k,i}_{22},\quad c_{k,i}=\frac{19l^{k,i}_{22}}{k^{1/4}},$ (46)
@@
In this subsection, we propose a completely parameter-free AGP algorithm for solving nonconvex-concave minimax problems. Similar to the idea of the PF-AGP-NSC Algorithm, we estimate the Lipschitz constants $L_{11}$, $L_{12}$ and $L_{22}$ iteratively, however we do not need to estimate the strongly concavity modulus $\mu$. At each iteration of the proposed algorithm, we aim to find a tuple $(l^{k,i}_{11},l^{k,i}_{12})$ by backtracking such that conditions (2.1), (C2). The proposed algorithm, denoted as PF-AGP-NC, is formally presented in Algorithm 2.
Please note that the existing SGDA-B algorithm (Xu et al., 2024) for solving nonconvex-concave minimax problems requires the realization of the diameter of the feasible set $\mathcal{Y}$ in order to construct the key auxiliary problem, while the proposed PF-AGP-NC algorithm is a completely parameter-free algorithm that does not require any prior parameters.
@@ 2.2.1 Complexity Analysis
In this subsection, we analyze the iteration complexity of Algorithm 2 for solving (P) under the general nonconvex-concave setting. Similar to the idea of the PF-AGP-NSC Algorithm, we estimate the Lipschitz constants $L_{11}$, $L_{12}$ and $L_{22}$ iteratively, however we do not need to estimate the strongly concavity modulus $\mu$. At each iteration of the proposed algorithm, we aim to find a tuple $(l^{k,i}_{11},l^{k,i}_{12},l^{k,i}_{22})$ by backtracking such that conditions (2.1), (C2) and the following condition (C5) are satisfied:
${C}_{5}^{k,i}=$ $(l_{22}^{k,i}+c_{k,i})\langle\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k}),y_{k,i}-y_{k}\rangle$
$+\|\nabla_{y}f_{k,i}(x_{k,i},y_{k,i})-\nabla_{y}f_{k,i}(x_{k,i},y_{k})\|^{2}\leq 0$ (C5)
The backtracking strategy (C5) is a key inequality that we will use in the following two lemmas to establish some important recursions of the PF-AGP-NC algorithm in the nonconvex-concave setting. This is also one of the main differences between proofs in the nonconvex-strongly concave setting and proofs in the nonconvex-concave setting.
@@
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. If $\forall k$, $\beta_{k}>l^{k}_{11}$ and  $\gamma_{k}\geq{{l}_{22}^{k}+c_{k}}$, then
@@
$f_{k}(x_{k+1},y_{k+1})-f_{k}(x_{k+1},y_{k})$
$\leq$ $\langle\nabla_{y}f_{k}(x_{k+1},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k}),y_{k+1}-y_{k}\rangle$
$+\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k+1}-y_{k}\rangle$
@@
$+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
$+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),y_{k}-y_{k-1}\rangle$
$+\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle,$ (49)
@@
Thirdly, by (C5) , we have that
@@
$\leq$ $-\frac{1}{{l}_{22}^{k-1}+c_{k-1}}\|\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1})\|^{2}.$ (52)
@@
$\gamma_{k-1}\langle y_{k}-y_{k-1},y_{k+1}-y_{k}\rangle$
$=$ $\frac{\gamma_{k-1}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}-\frac{\gamma_{k-1}}{2}\|v_{k+1}\|^{2}.$ (53)
By $\gamma_{k}\geq{{l}_{22}^{k}+c_{k}}$, plugging (50)-(53) into (49), and using the definition of $f_{k}(x_{k+1},y_{k+1})$ and $f_{k}(x_{k+1},y_{k})$, we obtain
@@
$\beta_{k}>l^{k}_{11},\quad\frac{\gamma_{k+1}}{c_{k+1}}-\frac{\gamma_{k}}{c_{k}}\leq\frac{1}{5},\quad{c_{k+1}}-{c_{k}}\leq{\gamma_{k+1}-\gamma_{k}},\quad\gamma_{k}\geq{{l}^{k}_{22}+c_{k}},$ (58)
@@
$F_{k+1}-F_{k}=$ $-\frac{19\gamma_{k}}{10}\|y_{k+1}-y_{k}\|^{2}-\left(\frac{\beta_{k}}{2}-\frac{(l^{k}_{12})^{2}}{2\gamma_{k-1}}-\frac{16(l^{k}_{12})^{2}\gamma_{k-1}}{c_{k-1}^{2}}\right)\|x_{k+1}-x_{k}\|^{2}$
@@
$\langle\nabla_{y}f_{k-1}(x_{k},y_{k-1})-\nabla_{y}f_{k}(x_{k+1},y_{k})+\gamma_{k}(y_{k+1}-y_{k})-\gamma_{k-1}(y_{k}-y_{k-1}),y_{k+1}-y_{k}\rangle$
$\leq 0,$
@@
$+\langle\nabla_{y}f_{k-1}(x_{k},y_{k})-\nabla_{y}f_{k-1}(x_{k},y_{k-1}),v_{k+1}\rangle$
@@
Using an argument similar to the proof of (50)-(53), by (C5), $f_{k}(x,y)$ is concave with respect to $y$ and the Cauchy-Schwarz inequality, we conclude from the above inequality that
@@
$-\frac{c_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2},$ (63)
for any $a_{k}>0$. Combining $\gamma_{k}\geq{{l}^{k}_{22}+c_{k}}$ and rearranging the terms in (63), we obtain
@@
$\quad\leq\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}+\left(\frac{a_{k}}{2}+\frac{\gamma_{k-1}}{2}-\frac{\gamma_{k}}{2}+\frac{c_{k}}{2}-\frac{c_{k-1}}{2}\right)\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k}\|^{2}$
$\quad\quad+\frac{(l^{k}_{12})^{2}}{2a_{k}}\|x_{k+1}-x_{k}\|^{2}-\frac{c_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}$
$\quad\leq\frac{\gamma_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}+\frac{a_{k}}{2}\|y_{k+1}-y_{k}\|^{2}+\frac{c_{k}-c_{k-1}}{2}\|y_{k}\|^{2}+\frac{(l^{k}_{12})^{2}}{2a_{k}}\|x_{k+1}-x_{k}\|^{2}$ (64)
$\quad\quad-\frac{c_{k-1}}{2}\|y_{k}-y_{k-1}\|^{2}.$ (65)
@@
Suppose that Assumptions 1 and 2 hold. Let $\{(x_{k},y_{k})\}$ be a sequence generated by Algorithm 2. If $\beta_{k}=\frac{l_{12}^{k}}{20l^{k-1}_{22}}+\frac{2(l_{12}^{k})^{2}\sqrt{k}}{l_{22}^{k-1}}$, $\gamma_{k}=20l^{k}_{22}$ , $c_{k}=\frac{19l^{k}_{22}}{k^{1/4}}$, then for any given $\varepsilon>0$, $k\geq 2$, we have
$T(\varepsilon)\leq\max\left(\left(\frac{\hat{\alpha}d_{2}d_{3}}{\varepsilon^{2}}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$
where $d_{2}=F_{2}-\underline{F}+339L_{22}\hat{\mathcal{D}}_{y}^{2}$, $d_{3}=\max\{\bar{d}_{1},\max\limits_{{k=2,\cdots,T(\varepsilon)}}\{\frac{19^{2}\cdot 20^{2}l_{22}^{k}l_{22}^{k-1}}{16(l_{12}^{k})^{2}\}}\}\}$, $\hat{\mathcal{D}}_{y}=\max\{\|y\|\,|y\in Y\}$ with $\underline{F}={\underline{f}}-639L_{22}\hat{\mathcal{D}}_{y}^{2}$, $\underline{f}=\min_{(x,y)\in X\times Y}f(x,y)$, $\hat{\alpha}=\max\limits_{{k=2,\cdots,T(\varepsilon)}}\{\frac{2^{8}\cdot 5^{2}(l^{k}_{12})^{2}}{19^{2}\cdot 20l_{22}^{k-1}}\}$ and $\bar{d}_{1}=\max\limits_{{k=2,\cdots,T(\varepsilon)}}\left\{4+\frac{19^{4}}{2^{15}\cdot 5^{4}}+\frac{19^{4}\cdot 20^{2}(l_{22}^{k})^{2}}{2^{15}\cdot 5^{4}\cdot(l_{12}^{k})^{2}}\right\}$.
@@
This observation, in view of Lemma 9, implies that
${\alpha_{k}}\left\|x_{k+1}-x_{k}\right\|^{2}+\frac{19\gamma_{k}}{10}\left\|y_{k+1}-y_{k}\right\|^{2}$
$\leq$ $F_{k}-F_{k+1}+8\left({\gamma_{k}}\frac{c_{k+1}}{c_{k}}-{\gamma_{k-1}}\frac{c_{k}}{c_{k-1}}\right)\|y_{k+1}\|^{2}+\frac{c_{k-1}-c_{k}}{2}\|y_{k+1}\|^{2}.$ (68)
@@
$\leq\max\limits_{k=2,3\cdots T(\varepsilon)}\{4+\frac{19^{4}}{2^{15}\cdot 5^{4}}+\frac{19^{4}\gamma^{2}_{k}}{2^{15}\cdot 5^{4}({l}_{12}^{k})^{2}}\}=\bar{d}_{1}.$ (73)
@@
Denote  $d_{k}^{(2)}=\frac{1}{\max\{\bar{d}_{1}\alpha_{k},{\frac{19}{20}\gamma_{k}}\}}$, $\tilde{c}_{k}=\frac{19}{20k^{1/4}}$, by multiplying $d^{(2)}_{k}$ on the both sides of (74), and using (68), we have
@@
$\sum_{k=2}^{\tilde{T}{(\varepsilon)}}d^{(2)}_{k}\left\|\nabla\tilde{G}_{k}\right\|^{2}$
$\leq$ $F_{2}-F_{\tilde{T}(\varepsilon)}+8\left(\frac{\gamma_{\tilde{T}(\varepsilon)-1}c_{\tilde{T}(\varepsilon)}}{c_{\tilde{T}(\varepsilon)-1}}-\frac{\gamma_{1}c_{2}}{c_{1}}\right)\hat{\mathcal{D}}_{y}^{2}+40L_{22}\left(\frac{\tilde{c}_{1}}{2}-\frac{\tilde{c}_{\tilde{T}(\varepsilon)}}{2}\right)\hat{\mathcal{D}}^{2}_{y}$
@@
We can see from the selection of $d_{3}$ that  $d_{3}\geq\max\{\bar{d}_{1},\frac{19\gamma_{k}}{20\alpha_{k}}\}$, which implies $d^{(2)}_{k}\geq\frac{1}{d_{3}\alpha_{k}}$, by multiplying $d_{3}$ on the both sides of (79), and combining the definition of $d_{2}$, we have $\sum_{k=2}^{\infty}\frac{1}{\alpha_{k}}\left\|\nabla\tilde{G}_{k}\right\|^{2}\leq d_{2}d_{3}$, which, by the definition of $\tilde{T}(\varepsilon)$, implies that
@@
Where $\bar{\alpha}={(k-1)^{1/2}}\hat{\alpha}$, by using the fact $\sum_{k=2}^{\tilde{T}{(\varepsilon)}}\frac{1}{\sqrt{k}}\geq\sqrt{\tilde{T}{(\varepsilon)}}-1$ and (80), we conclude  $\frac{\varepsilon^{2}}{4}\leq\frac{\hat{\alpha}d_{2}d_{3}}{\sqrt{\tilde{T}(\varepsilon)}-1}$ or equivalently,
$\tilde{T}(\varepsilon)\leq\left(\frac{\hat{\alpha}d_{2}d_{3}}{\varepsilon^{2}}+1\right)^{2}.$ (81)
@@
$\leq\max\left(\left(\left(\frac{\hat{\alpha}d_{2}d_{3}}{\varepsilon^{2}}+1\right)^{2}+1\right)^{2},\frac{76^{4}L_{22}^{4}\hat{\mathcal{D}}_{y}^{4}}{\varepsilon^{4}}\right),$ (82)